OpenStudy (anonymous):

Help? 1 step at a time :) derivative y=4log_8(5x+3)

OpenStudy (anonymous):

Im assuming that that means base 8 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. d/dx((4log(5x+3)/log(8)) Factor out constants: \[4(d/dx)((\log(5x+3)/\log(8))\] use chain rule: \[(4)(d/dx(5x+3)/5x+3)/\log(8)\] Differentiate the sum term by term and factor out the constants: \[4(5(d/dx(x)+d/dx(3))/(5x+3)\log(8)\] Derivative of x is 1: y'(x) = \[20/(5x+3)\log(8)\]